OpenStudy (anonymous):

Is this the right form: 2/((x^2+2)(x^2+1)) ---->(Ax+B)/(x^2+2)+(Cx+D)/(x^2+1)

OpenStudy (anonymous):

\[\frac{ 2 }{ (x ^{2}+2)(x ^{2}+1) } = \frac{ Ax+B }{ x ^{2}+2 }+\frac{ Cx+D }{ x ^{2}+1 }\]

hartnn (hartnn):

yes, it seems correct. do you want to integrate it ? or just partial fractions......?

OpenStudy (anonymous):

just partial fractions

hartnn (hartnn):

ok, continue.....u know the next steps?

OpenStudy (anonymous):

Yeah I just wasn't sure if that is the right form. Thanks!! :)

hartnn (hartnn):

welcome ^_^